Destroying the Trojan horse. Penicillin is hydrolyzed and thereby rendered inactive by penicillinase (also known as ?-lactamase), an enzyme present in some resistant bacteria. The mass of this enzyme in Staphylococcus aureus is 29.6 kd. The amount of penicillin hydrolyzed in 1 minute in a 10-ml solution containing 10-9 g of purified penicillinase was measured as a function of the concentration of penicillin. Assume that the concentration of penicillin does not change appreciably during the assay. (How do you use the Michaelis-Menten equation without Vmax or Km?)


1
Expert's answer
2022-01-19T12:37:02-0500

Conversion Required-

29.6kd = "4.915\times10^-20" g

number of molecule of enzyme present =

"\frac{10^-9}{4.915\times 10^-20}"

="2.03\times 10^-10" molecules

Total mass ="\frac{2.03 \times(10)^10}{6.023\times(10)^20 }"

=3.37"\times" 10"^-14"

Vmax = K2 "\times" n

6.8"\times" 10"^-10" mol/min = K2 "\times" 3.37"\times" 10"^-14" mol

K2 = 337 s"^-1"